OpenStudy (anonymous):

a train is moving 4.82 m/s when it begins to accelerate. after 24.8s it is moving 7.09 m/s. how far did it move in that time? HELP PLZ

OpenStudy (anonymous):

So what u do is rearrange the formula from your first question to get Distance=(velocity)*(time) where velocity =(7.09-4.82) and the time is (24.8-0) therefore the Distance=(7.09-4.82)*(24.8) Distance=56.296metres

OpenStudy (anonymous):

thank u but the answer is wrong :-/ i tried that already before and nothing :(

OpenStudy (anonymous):

whats the answer????

OpenStudy (anonymous):

i dont know. :/ im still stuck :(

OpenStudy (anonymous):

so how do u know if the answer above is wrong?

OpenStudy (anonymous):

because i am taking online classes and this is one of the question they have asked me to answer and i put this answer and it came out wrong. :/

OpenStudy (anonymous):

oh ok give me some time ill try work out this question and then help u out if that is alright

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u=4.82 m/s v=7.09 m/s t=24.8 s a=? s=? v=u+at 7.09=4.82+24.8a 24.8a=7.09-4.82=2.27 a=2.27/24.8 s=ut+1/2at^2 s=4.82*24.8+1/2 *2.27/24.8 *(24.8)^2 s=(4.82)*(24.8)+1/2 *2.27*24.8 s=24.8*(4.82+1.135) s=24.8*5.955 s=147.684 meters